# ToolACE-R: Tool Learning with Adaptive Self-Refinement

### v1

@@ Abstract
Tool learning, which allows Large Language Models (LLMs) to leverage external tools for solving complex user tasks, has emerged as a promising avenue for extending model capabilities. However, current approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model. In this paper, we propose ToolACE-R, a novel method that introduces adaptive self-refinement for tool invocations. Our approach features a model-aware iterative training procedure that progressively incorporates more training samples based on the model’s evolving capabilities. Additionally, it allows LLMs to iteratively refine their tool calls, optimizing performance without requiring external feedback. To further enhance computational efficiency, we integrate an adaptive mechanism when scaling the inference time, enabling the model to autonomously determine when to stop the refinement process. We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models, even without any refinement. Furthermore, its performance can be further improved efficiently through adaptive self-refinement. Our results demonstrate the effectiveness of the proposed method, which is compatible with base models of various sizes, offering a promising direction for more efficient tool learning.
ToolACE-R: Tool Learning with Adaptive Self-Refinement
Xingshan Zeng1, Weiwen Liu1, Xu Huang2, Zezhong Wang3, Lingzhi Wang4, Liangyou Li1, Yasheng Wang1, Lifeng Shang1, Xin Jiang1, Ruiming Tang1, Qun Liu1 1Huawei Noah’s Ark Lab 2University of Science and Technology of China 3The Chinese University of Hong Kong 4Harbin Institute of Technology, Shenzhen zeng.xingshan@huawei.com, liuweiwen8@huawei.com, xuhuangcs@mail.ustc.edu.cn
@@ 1 Introduction
Tool learning, which enables Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention. With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans Qu et al. (2024b). While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools Qin et al. (2023); Patil et al. (2023); Liu et al. (2024b, c).
@@
Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models Liu et al. (2024b, c); Wang et al. (2024b); Abdelaziz et al. (2024). However, data synthesized by advanced models from different sources can lead to compatibility issues. Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations Ren et al. (2024); Kang et al. (2024). As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
@@
Additionally, the potential of maximizing the model’s own capabilities remains under exploration in the field of tool learning. It has been shown that a model’s reasoning performance can be improved by scaling test-time computation Brown et al. (2024); Snell et al. (2024). One promising approach involves iteratively refining the model’s output multiple times and then selecting the majority answer Snell et al. (2024). However, user queries vary in complexity, ranging from simple to highly complex. Scaling test-time computation for overly simple and straightforward questions is inefficient, as models can often answer them correctly without additional refinement. Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of effort required for each question.
@@
To address the challenges outlined above, we propose ToolACE-R, a method that enables adaptive self-refinement of tool invocations through a customized iterative training procedure. Specifically, we introduce a model-aware iterative training approach that facilitates LLMs in learning self-refinement for tool invocation in an iterative fashion, enhanced by model-based data selection. Additionally, by constructing refining data samples that preserve identical cases, LLMs are more likely to learn appropriate stopping criteria during the iterative process. Building on this training procedure, we further integrate adaptive self-refinement into iterative inference for inference time scaling, allowing models to autonomously determine when to halt the iteration, thereby improving scaling efficiency.
@@
We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) Yan et al. (2024) and API-Bank Li et al. (2023), to assess the effectiveness and efficiency of ToolACE-R.The experimental results show that ToolACE-R achieves competitive performance compared to advanced API-based models like GPT-4o, and can be further enhanced efficiently through adaptive self-refinement. Additional analyses highlight the impact of the proposed modules and examine the generalizability of the method across different model backbones and sizes.
@@
  • We propose a model-aware iterative training procedure, enabling LLMs learning adaptive self-refinement for tool learning, which maximizes the model potential without any feedback.
  • We integrate adaptive self-refinement into iterative inference, enabling models to autonomously decide when to halt the refinement process and allowing for more efficient use of computational resources when scaling test-time compute.
@@ 2 Related Work
Integrating external tools can enhance LLMs’ capabilities, enabling more specialized, precise, and reliable problem-solving for users (Qin et al., 2023). Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches. Tuning-free methods employ prompting techniques, such as in-context tool-use examples, without additional training Mialon et al. (2023); Hsieh et al. (2023); Ruan et al. (2023). A notable example method is ReAct Yao et al. (2023), which can alternate between reasoning and action for complex tasks. Recent work Qu et al. (2024a) improves tool understanding by iteratively refining documentation. Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention Qin et al. (2023); Schick et al. (2023); Patil et al. (2023); Tang et al. (2023); Abdelaziz et al. (2024). Due to data scarcity, a large amount of them focus on data synthesis for improved generation procedures Liu et al. (2024b, c); Wang et al. (2024b), often overlooking the model’s inherent potential. While some others integrate tool feedback for refinement Du et al. (2024); Wang et al. (2024a), our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
@@
Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs Albalak et al. (2024). When the base model is sufficiently powerful, a small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of general data Zhou et al. (2023); Liu et al. (2023). While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection Du et al. (2023); Li et al. (2024). This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance Ren et al. (2024). Our method aligns with this perspective, proposing the use of pass@k metric to identify and select suitable training samples based on the model to be trained.
@@
Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback Madaan et al. (2023); Weng et al. (2023) or external feedback Qu et al. (2024c); Xu et al. (2024). Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously Huang et al. (2024), which means that scaling test-time compute relying on iterative refinement still often depend on post-processing techniques, such as self-consistency Snell et al. (2024). In this work, we propose a simple yet effective method that enables models to adaptively self-refine their outputs, thereby improving scaling efficiency.
Figure 1: The overview of ToolACE-R, which includes two components, for training and inference, respectively.
@@ 3.1 Problem Formulation
where $t_{j}$ and $a_{j}$ represent the $j$-th invoked tool and its corresponding argument, respectively, with $1\leq j\leq m$ and $m$ being the total number of tool invocations needed. The function $f_{\theta}(\cdot)$ denotes the generation process of the LLM with parameters $\theta$.
The training sample for tool learning is typically formalized as a triplet consisting of a user query, candidate tools, and the ground-truth tool invocation: $\{\langle q,T\rangle,A\}$, where $q$ and $T$ serve as input and $A$ should be output.
@@ 3.2 Model-Aware Iterative Training
Our model training starts with an instructed base model, ensuring that it has basic instruction following capability for our model-aware module. For training, we first gather a set of off-the-shelf training samples. Instead of using them directly, we apply a selection criterion to curate the data, which is then combined with generated self-refinement data. This curated and refined data is integrated into an iterative training process.
Data Selection Criterion.
Inspired by self-improvement techniques in existing LLM reasoning research Zelikman et al. (2022), where LLMs are trained using their own outputs, we introduce a pass@k criterion to select the most suitable training samples for each base model – those the model can answer (almost) correctly within at most $k$ tries. This approach may seem counterintuitive, as one might expect models to focus on learning from samples they struggle with. However, recent research Ren et al. (2024) has highlighted that models are more likely to perform self-alignment rather than learning new information during the fine-tuning stage. Based on this insight, we adopt this pass@k metric for data selection.
Specifically, for each sample $\{\langle q,T\rangle,A\}$, the base model to be trained first generate the invocation up to $k$ times. If any of the responses approaches the groundtruth $A$, we retain it as an official training sample:
$R(\langle q,T\rangle,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (2)
Here $R(\langle q,T\rangle,A)=1$ indicates the training sample will be retained. $\theta_{0}$ represents the base model before training and $f_{\theta_{0}}(\cdot)_{j}$ means the $j$-th generation with the base model.
A key aspect is how to define when a response "approaches" $A$. In our preliminary experiments, we found that if we restrict them to be exactly equal, the resulting training set still has an effect but limits the potential for performance improvement (see Table 4). However, by loosening this constraint, more suitable training samples are retained and the model’s potential can be more effectively harnessed. Therefore, we use the concept of "approaching" in this context. For any two invocations $A$ and $A^{\prime}$, we define $A\approx A^{\prime}$ if the used tools are exactly the same and each invocation only contains at most 1 different parameter values.
Self-Refine Data Construction.
@@
To establish the model’s self-refinement capability, we additionally construct self-refine data based on the training samples. Specifically, we create a self-refine sample as a multi-turn interaction, where the first turn is the initial response from the model, and the second turn includes a refinement prompt from the user followed by the model’s refined answer: $\{\langle q,T\rangle,A_{1},r,A_{2}\}$. We apply the same selection criterion as described above, retaining the self-refinement training sample only when $A_{2}\approx A$:
$R(\langle q,T\rangle,A_{1},r,A)=\begin{cases}1&\text{if }\exists\,j\in\{1,2,..,k\}\,\text{s.t. }\\ &f_{\theta_{0}}(\langle q,T\rangle,A_{1},r)_{j}\approx A,\\ 0&\text{otherwise}.\end{cases}$ (3)
Here, we replace $A_{2}$ with $A$ to create the refined training sample: $\{\langle q,T\rangle,A_{1},r,A\}$.
@@
Iterative Training.
Given the selected training samples (including self-refine samples), we apply supervised fine-tuning to the base model for one epoch. This constitutes one iteration. In subsequent iterations, we treat the trained model as the new base model, re-construct the self-refine samples, and re-select those that are suitable for fine-tuning the updated model. The model is then further trained using the newly selected samples.
Ideally, with each iteration, the model’s performance improves, leading to an increase in the number of selected samples. The iteration process continues until the number of selected samples no longer increases, indicating that the model has saturated on this set of training samples. The resulting model will then be our final ToolACE-R model.

### v2

# ToolACE-R: Model-aware Iterative Training and Adaptive Refinement for Tool Learning

@@ Abstract
Tool learning, which allows Large Language Models (LLMs) to leverage external tools for solving complex user tasks, has emerged as a promising avenue for extending model capabilities. However, existing approaches primarily focus on data synthesis for fine-tuning LLMs to invoke tools effectively, largely ignoring how to fully stimulate the potential of the model. In this paper, we propose ToolACE-R, a novel framework that includes both model-aware iterative training and adaptive refinement for tool learning. ToolACE-R features a model-aware iterative training procedure that progressively adjust training samples based on the model’s evolving capabilities to maximize its potential. Additionally, it incorporates self-refinement training corpus which emphasizes LLM’s ability to iteratively refine their tool calls, optimizing performance without requiring external feedback. Furthermore, we introduce adaptive self-refinement mechanism for efficient test-time scaling, where the trained model can autonomously determine when to stop the process based on iterative self-refinement. We conduct extensive experiments across several benchmark datasets, showing that ToolACE-R achieves competitive performance compared to advanced API-based models. The performance of tool invocation can be further improved efficiently through adaptive self-refinement. These results highlight the effectiveness and generalizability of ToolACE-R, offering a promising direction for more efficient and scalable tool learning.
@@ 1 Introduction
Tool learning, which enables Large Language Models (LLMs) to leverage external tools to address complex user requirements, has gained increasing attention. With tool integration, LLMs can access up-to-date information, perform intricate computations, and utilize third-party services, significantly expanding their capabilities beyond simple natural language communication with humans [27]. While tool invocation requires the LLMs to demonstrate strong understanding, reasoning, and instruction-following skills, customized fine-tuning is currently the dominant approach for enabling models to call external tools [19, 20, 23, 25].
@@
Due to the limited availability of high-quality data, existing research has primarily focused on developing effective and efficient methods for data synthesis using advanced models [1, 19, 20, 35]. However, data synthesized by advanced models from different sources can lead to compatibility issues. Specifically, when synthesized samples exceed the model’s current knowledge, they may undermine the model’s performance or lead to hallucinations [12, 29]. As a result, determining which training data samples are appropriate for a given model remains an unresolved challenge.
@@
Additionally, the potential for maximizing a model’s intrinsic capabilities remains underexplored in the field of tool learning. On one hand, prior work has seldom investigated how to fully exploit existing training data through techniques such as data augmentation. On the other hand, scaling test-time computation – despite its demonstrated effectiveness in enhancing LLM reasoning [4, 32] – has received limited attention in tool learning scenarios. A promising approach involves iteratively refining the model’s outputs and selecting the majority answer [32]. However, user queries vary widely in complexity, ranging from simple to highly intricate. Scaling test-time computation indiscriminately is inefficient, as models can often correctly answer straightforward questions without the need for additional refinement. Therefore, an adaptive strategy for scaling test-time computation is needed to dynamically determine the appropriate level of computational effort required for each query.
@@
To address the challenges outlined above, we propose ToolACE-R, a novel pipeline for stimulating model potential in tool learning, encompassing both training and inference procedures. Specifically, we propose a model-aware iterative training framework that enables LLMs to learn tool invocation in alignment with their evolving capabilities. This is achieved through a novel model-aware difficulty metric, guiding the training process in an iterative manner. Additionally, a simple yet effective self-refinement data incorporation strategy is employed as a form of data augmentation. Furthermore, by preserving refinement samples that include identical cases, LLMs are encouraged to learn appropriate stopping criteria during the iterative process. Building upon this training procedure, we integrate adaptive self-refinement into iterative inference, allowing models to autonomously determine when to halt refinement, thereby improving inference-time efficiency.
@@
We have conducted extensive experiments on several representative tool-calling benchmarks, such as the Berkeley Function Call Leaderboard (BFCL) [38] and API-Bank [15], to assess the effectiveness and efficiency of ToolACE-R. Experimental results demonstrate that ToolACE-R achieves competitive performance compared to advanced API-based models such as GPT-4o, and can be further enhanced through adaptive self-refinement. Additional analyses highlight the contributions of the proposed modules and examine the generalizability of our method across different model backbones and sizes.
@@
  • We propose a model-aware iterative training procedure, augmented with self-refinement corpus construction, to maximize the potential of LLMs for tool learning.
  • Building upon the training process, we integrate adaptive self-refinement into iterative inference, enabling models to autonomously decide when to halt the refinement process and allowing for more efficient use of computational resources when scaling test-time compute.
@@ 2 Related Work
Tool learning methods generally fall into two categories, tuning-free and tuning-based approaches. Tuning-free methods employ prompting techniques without additional training [22, 8, 30]. A notable example method is ReAct [40], which can alternate between reasoning and action for complex tasks. Recent work [26] improves tool understanding by iteratively refining documentation. Tuning-based methods, which directly enhance tool calling ability via specialized fine-tuning, have recently gained much more attention [25, 31, 23, 33, 1]. A large amount of them focus on data synthesis for improved generation procedures [19, 20, 35], often overlooking the model’s inherent potential. While some others integrate tool feedback for refinement [7, 34], our approach is more streamlined, directly refining outputs in an iterative manner without external or any textual feedback to maximize model efficiency.
@@
Research indicates that selecting high-quality training samples is essential for fine-tuning LLMs [3]. A small set of high-quality data can effectively harness the model’s potential, rather than relying on large quantities of data [42, 18]. While earlier works primarily emphasize general data quality aspects, such as diversity and complexity, recent studies advocate for model-specific data selection [6, 14]. This kind of approach is based on the observation that data distributions that significantly deviate from the base model’s are challenging for the model to learn from and may even degrade the performance [29]. Our method aligns with this perspective, defining a metric with passrate@k to identify and select suitable training samples based on the model to be trained.
@@
Previous research has demonstrated that LLMs can refine their own outputs through either self-feedback [21, 36] or external feedback [28, 37]. Nevertheless, it remains challenging for LLMs to assess the correctness of their refined output autonomously [10], which means that scaling test-time compute with iterative refinement still often depend on post-processing techniques, such as self-consistency [32]. In this work, we propose a simple yet effective method that enables models to adaptively self-refine their outputs, thereby improving scaling efficiency.
Figure 1: The overview of ToolACE-R, including two components, for training and inference, respectively.
@@ 3.1 Problem Formulation
where $t_{j}$ and $a_{j}$ represent the $j$-th invoked tool and its corresponding argument, respectively, with $1\leq j\leq m$ and $m$ being the total number of tool invocations needed. The function $f_{\theta}(\cdot)$ denotes the generation process of the LLM with parameters $\theta$. For each $a_{j}$, it might include several parameters and the corresponding values, denoted as $a_{j}=[p_{1}:v_{1},...,p_{i}:v_{i},$...], where $p_{i}$ is the parameter name, and $v_{i}$ is the corresponding value.
The training samples for tool learning typically leverage $q$ and $T$ as context, and $A$ as output (the ground-truth tool invocation), denoted as $\{\langle q,T\rangle,A\}$.
@@ 3.2 Model-Aware Iterative Training
We begin model training with an instruction-tuned base model to ensure it possesses fundamental instruction-following capabilities required by our model-aware module. For the training process, we first collect a set of off-the-shelf samples and augment them via self-refinement. Rather than using these samples directly, we apply a selection criterion grounded in our defined notion of model-aware difficulty to curate the data. This curated dataset is then used to train the model and is iteratively re-curated based on new trained model throughout the training process.
Model-Aware Difficulty.
To guide tool learning with finer granularity, we introduce a model-aware difficulty metric that quantifies the inherent challenge a specific training sample poses to a given model. Rather than relying solely on static heuristics or dataset-level statistics, our definition measures how well the model can approximate the desired tool invocation behavior before training, thereby capturing a performance-aware notion of difficulty.
Let a training sample be denoted as $\{\langle q,T\rangle,A\}$, where $A=[t_{1}(a_{1}),\dots,t_{m}(a_{m})]$ is the reference tool invocation sequence. Given a base model $\theta_{0}$, we obtain the model’s prediction as $A^{\prime}=[t^{\prime}_{1}(a^{\prime}_{1}),\dots,t^{\prime}_{m^{\prime}}(a^{\prime}_{m^{\prime}})]=f_{\theta_{0}}(\langle q,T\rangle)$. To assess the similarity between the predicted and reference sequences, we adopt Jaccard-style overlap to define a parameter-level alignment score between individual tool calls $t_{i}(a_{i})$ and $t^{\prime}_{j}(a^{\prime}_{j})$ as follows:
$S_{ij}=\begin{cases}\frac{|a_{i}\cap a^{\prime}_{j}|}{|a_{i}\cup a^{\prime}_{j}|}&\text{if }t_{i}=t^{\prime}_{j}\\ 0&\text{otherwise }\end{cases}$ (2)
Here, $|a_{i}\cap a^{\prime}_{j}|$ denotes the number of matching arguments between $a_{i}$ and $a^{\prime}_{j}$, i.e., parameters with both identical names and values. $|a_{i}\cup a^{\prime}_{j}|$ represents the total count of distinct arguments across both calls. Parameters with the same name but differing values are counted separately to penalize incorrect assignments.
We aggregate all pairwise scores into a similarity matrix $S\in\mathbb{R}^{m\times m^{\prime}}$. To compute the best structural alignment between $A$ and $A^{\prime}$, we frame the task as a bipartite matching problem over the index sets $\mathcal{I}={1,\dots,m}$ and $\mathcal{J}={1,\dots,m^{\prime}}$, and apply the Hungarian algorithm [13] to identify the optimal matching $\mathcal{M}$ that maximizes the total alignment score:
$\mathcal{M}={\arg\max}_{\mathcal{M}\in\text{Matchings}(\mathcal{I},\mathcal{J})}\sum_{(i,j)\in\mathcal{M}}S_{ij}$ (3)
where $\text{Matchings}(\mathcal{I},\mathcal{J})$ denotes the set of all one-to-one matchings between $\mathcal{I}$ and $\mathcal{J}$, i.e., subsets of $\mathcal{I}\times\mathcal{J}$ in which each element in $\mathcal{I}$ and $\mathcal{J}$ appears in at most one pair. Using this optimal alignment, we define the overlap score between $A$ and $A^{\prime}$ as:
$O(A,A^{\prime})=\frac{\sum_{(i,j)\in\mathcal{M}}S_{ij}}{\max(m,m^{\prime})}$ (4)
This score reflects the degree to which the model reproduces the reference tool invocation, normalized by the longer of the two sequences to account for missing or superfluous tool calls.
To incorporate model uncertainty and account for variance in its outputs, we evaluate the model over $k$ independent attempts, resulting in outputs $A^{\prime(1)},\dots,A^{\prime(k)}$}. The model-aware difficulty of a sample is then defined as:
$D(\langle q,T\rangle,A)=1-\frac{1}{k}\sum_{l=1}^{k}O(A,A^{\prime(l)})$ (5)
This difficulty score $D\in[0,1]$ quantifies the average divergence between the model’s output and the reference across multiple attempts. A lower score indicates that the model consistently produces accurate or nearly accurate tool calls, while a higher score suggests that the model struggles with the sample, failing to replicate even partial correctness. For instance, $D=0$ implies perfect performance across all outputs, whereas $D=1$ reflects complete failure in reproducing any reference calls.
This model-aware difficulty serves as a valuable signal for iterative training, enabling more effective model development in tool-use scenarios.
Data Augmentation with Self-Refinement.
@@
To augment the training corpus and establish the model’s self-refinement capability, we additionally construct self-refinement data based on the original training samples. Specifically, we create a self-refinement sample as a multi-turn interaction, where the first turn is the initial response from the model, and the second turn includes a refinement prompt from the user followed by the refined answer (typically the ground-truth tool invocation $A$): $\{\langle q,T\rangle,A_{1},r,A\}$. Here, the model to be trained will first produce its output tool invocation $A_{1}$. We then concatenate a refinement requirement prompt $r$ (a generalized sentence like “Please refine your answer…”) and the final ground-truth answer $A$.
@@
Each self-refinement sample can also be estimated its difficulty using Eq. 5, by replacing $\{\langle q,T\rangle\}$ with $\{\langle q,T\rangle,A_{1},r\}$ as context. In this way, all the training samples including the constructed self-refinement samples can be estimated and labeled with a difficulty score.
Model-Aware Iterative Data Selection and Training.
We hypothesize that the most suitable training samples are those within the model’s current capability, either too difficult or too simple samples are ineffective for model training. To that end, we empirically set two thresholds $\alpha$ and $\beta$, where $\alpha<\beta$. For any training sample $\{C,A\}$ (where $C$ indicates the context, $\langle q,T\rangle$ or $\{\langle q,T\rangle,A_{1},r\}$ for self-refinement data), if its estimated difficulty score $\alpha<D(C,A)<\beta$, we retain it as an effective training sample and add to the actual training set $\mathcal{D}_{0}$.
Given the selected training set $\mathcal{D}_{0}$, we apply supervised fine-tuning to the base model $\theta_{0}$ for one epoch, producing $\theta_{1}$. This constitutes the first iteration. In subsequent iterations, we treat the trained model as the new base model, re-construct the self-refinement samples, re-evaluate the difficulty, and re-select samples with the same criterion for fine-tuning the updated model.
Ideally, with each iteration, the model’s performance improves, leading to an update of data difficulty and so new training samples might be included matching model’s current capability. The iteration will continue until the performance improvement is marginal, indicating that the model has saturated on this set of training samples. The resulting model will then be our final ToolACE-R model.